A bridge whose total weight is 4500 N is 20 m long and supported at each end. Find the forces exerted at each end when a
1600 N tractor is located 5 m from the left end.
1
Expert's answer
2019-06-21T08:24:16-0400

From the equilibrium:

"W+F=R_L+R_R"

Take the moments about left end:


"R_R (20)=W(10)+F(5)"

"R_R (20)=4500(10)+1600(5)"

"R_R=2650\ N"

Thus,


"4500+1600=R_L+2650"

"R_L=3450\ N"